### v1

# Ten (or more!) reasons to register your software with the Astrophysics Source Code Library

Abstract
This presentation covered the benefits of registering astronomy research software with the Astrophysics Source Code Library (ASCL, ascl.net), a free online registry for software used in astronomy research. Indexed by ADS and Clarivate’s Web of Science, the ASCL currently contains over 3600 codes, and its entries have been cited over 17,000 times. Registering your code with the ASCL is easy with our online submissions system. Making your software available for examination shows confidence in your research and makes your research more transparent, reproducible, and falsifiable. ASCL registration allows your software to be cited on its own merits and provides a citation method that is trackable and accepted by all astronomy journals, and by journals such as Science and Nature. Adding your code to the ASCL also allows others to find your code more easily, as it can then be found not only in the ASCL itself, but also in ADS, Web of Science, and Google Scholar.
1Astrophysics Source Code Library (ASCL), Houghton, MI, USA; aallen@ascl.net, kduprie@stsci.edu
2Astronomy Department, University of Maryland, College Park, MD, USA
3Space Telescope Science Institute, Baltimore, MD, USA; kduprie@stsci.edu
1 Introduction
The Astrophysics Source Code Library [^1] https://ascl.net (ASCL) is a free online repository and registry of source codes used in astronomy research that have appeared in or been submitted to peer-reviewed publications. Established at Michigan Technological University in 1999 by Robert Nemiroff and John Wallin (Nemiroff & Wallin 1999), its editors examine peer-reviewed papers for computational methods and create entries for the found software to the library. The ASCL also accepts submitted entries that meet its criteria. [^2] https://ascl.net/home/getwp/3593 With over 3,500 entries, the ASCL covers a significant number of the astrophysics source codes used in peer-reviewed studies.
This poster presentation encouraged researchers to register their software with the ASCL by sharing twelve reasons – ten (or more!) – they should consider doing so.
2 The Ten (or more!) Reasons
Though the ASCL was started primarily to share software and make research more transparent, it offers additional benefits to software authors and the astronomy community.
2.1 Get a unique identifier: an ASCL ID
Software registered by the ASCL is assigned a unique identifier. This ASCL ID has the form ascl:yymm.xxx, where yy is the year, mm is the month, and xxx is an incremental number; each new month’s incremental numbers start at .001.
An ASCL ID is assigned when a code is accepted by the ASCL. This unique identifier is recognized by NASA’s Astrophysics Data System (ADS), the resolution service Identifiers.org, [^3] https://registry.identifiers.org/registry/ascl the Web of Science Data Citation Index, [^4] https://clarivate.com/academia-government/scientific-and-academic-research/research-discovery-and-referencing/web-of-science/data-citation-index/ Science, Nature, and other journals, and by publishers including AAS Publications and Elsevier.
2.2 It’s free
No fee; it’s free! Free to submit code, free to find code. And it’s free to download all of our entries in JSON, too, which you can do with the URL https://ascl.net/code/json.
2.3 Makes your code citable on its own merit
Because the ASCL ID is unique to a code, it can be, and often is, used to cite computational methods in journal articles, proceedings, and in theses. The ability to cite software is just as important as the ability to cite a scientific paper; the software citation provides credit to the author(s) of the code and improves the transparency and reproducibility of the research for which it was used.
When a journal article has been written about a piece of software, that software paper is often used to cite the code. However, the FORCE11 Software Citation Guidelines (Smith et al. 2016), which many publishers support, state that “Software citations should be accorded the same importance in the scholarly record as citations of other research products, such as publications and data”. The FORCE11 guidelines also recognize the value of software papers, and say that “a request from the software authors to cite a paper should typically be respected, and the paper cited in addition to the software.” (Emphasis ours.) The ASCL strongly supports this, and when a code author has indicated how their code should be cited, that specification is included in ASCL entries in the “Preferred citation method” field.
Many codes do not have a dedicated software paper, instead having been written to support research. Although the code author will sometimes request that the research paper be cited if the software is used, more often than not there is no clear way to cite the code. The ASCL can help with that, as the ASCL ID can be used for the citation. [^5] https://ascl.net/wordpress/about-ascl/citing-ascl-code-entries/
2.4 Indexed by ADS and Web of Science/Allows others to find your code easily
ASCL entries are indexed – ingested – by ADS and Web of Science. This not only broadens opportunity to learn about software registered by the ASCL from these indexing services, but because ADS’s holdings also appear in Google Scholar, ASCL entries are found there as well. For example, the software PyBDSF (Python Blob Detection and Source Finder) has an ASCL entry, [^6] https://ascl.net/1502.007 appears in ADS, [^7] https://ui.adsabs.harvard.edu/abs/2015ascl.soft02007M and also in Google Scholar, [^8] https://scholar.google.com/scholar?hl=en&as_sdt=0%2C47&q=PyBDSF making information about this package widely available not only to astronomers, but to those in other fields as well.
2.5 Citations to ASCL entries are trackable by ADS
Citations are trackable and counted by an indexer only if they appear in an article’s bibliography and the indexer doing the tracking can match bibliographic information to resources it contains. Because ADS ingests ASCL entries, ADS tracks and counts the citations to ASCL IDs that are found in the journals and other resources it indexes. As of the writing of this paper, ADS shows more than 17,000 citations to ASCL entries, as can be seen in Figure 1, including one code with over 800 citations. More than 41% of the codes in ASCL show citations in ADS by ASCL ID.
2.6 Citations by ASCL ID accepted by over 250 journals including all major astronomy journals
Citations by ASCL ID appear in more than 250 journals indexed by ADS, the top three of which are The Astrophysical Journal, Monthly Notices of the Royal Astronomical Society, and Astronomy & Astrophysics. There are, of course, more citations for codes listed in ASCL than indicated above, as not all codes are cited by ASCL ID.
Figure 1.: Citation count to ASCL entries as captured by ADS, retrieved 2024/12/15
ASCL IDs are used for citation even in journals outside astronomy (Shamir 2024), as is seen for the UDAT code (Shamir 2017). These citations may be captured by indexers such as Web of Science and Google Scholar.
2.7 Associated links from ADS entries for papers to ASCL entries
ADS uses links sent to it by the ASCL (and others) to link articles to software, software to data, data to articles, and so on, as shown in Figure 2. These links reveal the different objects necessary to convey the full picture of a research project, making it easier to find them.
Figure 2.: Related materials for the PyRADS software, as shown by ADS on 2024/12/15.
2.8 Easy submissions form
ASCL editors encourage software authors to submit their codes. Really! Please submit your software! ASCL has an easy-to-use online submissions form [^9] https://ascl.net/code/submit that is always available, with helpful information provided throughout the form. Submitted entries are displayed immediately for public viewing, and those that meet the resource’s criteria [^10] https://ascl.net/home/getwp/3593 are assigned an ASCL ID.
2.9 Largest curated indexed resource for astronomy research codes in existence, with 3,500+ codes
ASCL entries are curated; links to code sites are checked at least twice a week, with results posted on the ASCL’s Dashboard, [^11] https://ascl.net/dashboard and records are examined for overall health regularly. The library covers all aspects of computational astrophysics. The resource has a full-text search and can serve as a discovery tool for software for one’s own research or by students to learn how common astronomical problems are solved computationally. Codes registered by the ASCL can also serve as benchmarks for one’s own solutions to these problems.
2.10 Shows confidence in your research
Sharing your software, making your source code open for examination, and registering it so others can find it shows your integrity and adherence to the scientific method.
2.11 Makes astronomy research more falsifiable, transparent, reproducible … better!
Science is a “show me” endeavor. Computational methods are methods, and science demands that research methods be revealed. This allows others not only to see how your research was conducted, but also enables others to verify your methods and build upon them.
3 Conclusion
A major goal of the ASCL is to strengthen and support published research findings by making the computational methods used in this research discoverable and available for examination. Registering your software makes research more transparent and also provides other benefits, including:
  • making software citable in a trackable way, thereby providing a way for software authors to accrue citations for their computational contributions to the discipline;
  • making it possible to easily find software that one may want to use, thereby making the field more efficient; and,
  • finding software one may want to examine to learn from, which is valuable not only for early career researchers, but also for experienced researchers looking to upgrade their skills and for teaching computational methods to students.
Acknowledgments
The ASCL thanks the Heidelberg Institute of Theoretical Studies, University of Maryland College Park, and Michigan Technological University for their support, and research software authors everywhere for writing the codes that enable research.